Find the margin of error for the given values of​ c, sigma​, and n. cequals0.90​, sigmaequals2.8​, nequals36
Answer #1

Solution :

Given that,

\sigma = 2.8

n = 36

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

Margin of error = E = Z\alpha/2* (\sigma\sqrt/n)

= 1.645 * (2.8 / \sqrt 36 )

= 0.8

Margin of error = E = 0.8